# ViTextVQA: A Large-Scale Visual Question Answering Dataset for Evaluating Vietnamese Text Comprehension in Images

### v2

@@ 7.2 Vietnamese Word Segmentation
Table 9: Statistic of word segmented in OCR text with different arrangements type.
Arrangement Type Min Word Segmented Max Word Segmented Avg Word Segmented Avg EM (%) Avg F1-Score (%) Not sort 0 11 0.56 17.81 39.46 Confidence score sort 0 10 0.56 16.39 37.38 Top-left to bottom-right sort 0 14 1.06 19.28 41.73
@@ 8 Conclusion and Future Work
In this article, we presented a new dataset for the VQA task in Vietnamese to the global research community, especially Vietnam. Specifically, our ViTextVQA dataset includes more than 16,000 images and more than 50,000 question-answer pairs focusing on exploiting text appearing in images. We believe that the effort to create our ViTextVQA dataset will contribute to improving the diversity and enriching resources for the Vietnamese language.
@@ Acknowledgements
Appendix
Example Results of Models with OCR Arrangements
Figure 25: Example results of models with not sort.
Figure 26: Example results of models with confidence score sort.
Figure 27: Example results of models with top-left to bottom-right sort.

### v3

@@ 7.2 Vietnamese Word Segmentation
Table 12: Statistics of word segmentation in OCR text with different arrangement types.
Arrangement Method Min Max Avg Avg EM (%) Avg F1-score (%)
Not sorted 0 11 0.56 18.45 40.62
Confidence score sorted 0 10 0.56 17.30 39.04
Top-left to bottom-right sorted 0 14 1.06 19.83 42.81
@@ 8 Conclusion and Future Work
In this article, we presented a new dataset for the VQA task in Vietnamese to the global research community, especially Vietnam. Specifically, our ViTextVQA dataset includes more than 16,000 images and more than 50,000 question-answer pairs focusing on exploiting . We believe that the effort to create our ViTextVQA dataset will contribute to improving the diversity and enriching resources for the Vietnamese language.
@@ Acknowledgements
Author Contributions Statement
Quan Van Nguyen: Conceptualization; Formal analysis; Investigation; Methodology; Validation; Visualization; Writing - original draft.
Dan Quang Tran: Conceptualization; Data curation; Formal analysis; Investigation; Validation; Visualization; Writing - original draft.
Huy Quang Pham: Conceptualization; Data curation; Investigation; Methodology; Writing - original draft.
Thang Kien-Bao Nguyen: Conceptualization; Data curation; Investigation; Methodology; Writing - original draft.
Nghia Hieu Nguyen: Conceptualization; Data curation; Investigation; Methodology; Writing - review & editing.
Kiet Van Nguyen: Conceptualization; Formal analysis; Investigation; Methodology; Validation; Supervision; Writing - review & editing.
Ngan Luu-Thuy Nguyen: Conceptualization; Formal analysis; Investigation; Methodology; Validation; Supervision; Writing - review & editing.